164.0 Free Cash Flows Rhodes Corporation: Income Statements for Year Ending December 31 (Millions of Dollars) 2016 2015 Salesa. What is the net operating profit after taxes (NOPAT) for 2016? Enter your answer in millions. For example, an answer of $1Quantitative Problem: Rosnan Industries 2013 and 2012 balance sheets and income statements are shown below. Balance Sheets:The balance in the firms cash and equivalents account is needed for operations and is not considered excess cash. What is
Answer #1

RHODES CORPORATION:

1) NOPAT (net operating profit after tax) - 2016:

= EBIT (1 - tax rate)

= 2141 x (1 - 0.4) = $1284.6 million.

2) NET OPERATING WORKING CAPITAL:

= operating current assets - operating current liabs.

= cash + acc. receivables + inventory - (account payables + accrued expenses)

2016 2015
Operating Current Assets 3,097 2,389
Operating Current Liabs. 1,335 1,105
NET OPERATING WC 1,762 1,284

3) NET OPERATING CAPITAL:

= Net Operating Working Capital + Non-current Operating Assets

PARTICULARS 2016 2015
Operating Current Assets 3,097 2,389
Operating Current Liabs. 1,335 1,105
NET OPERATING WC (A) 1,762 1,284
Net PPE 1,964 1,785
NON-CURRENT OPERATING ASSETS (B) 1,964 1,785
NET OPERATING CAPITAL (A+B) 3,726 3,069

4) FREE CASH FLOW (2016):

= Operating Cash Flow − Capital Expenditures

= Net income + non-cash expenses - increase in WC - Capital expenditures

= 1164 + 196 + 638 + 375

= $2,373 milion.

5) ROIC (return on invested capital) - 2016:

= Net operating profit after tax / invested capital x 100

= 1284.6 x 100 / (1870 + 1716)

= (1284.6 x 100) / 3586

= 35.8%

ROSNAN INDUSTRIES:

1) NET OPERATING WORKING CAPITAL:

= operating current assets - operating current liabs.

= cash + acc. receivables + inventory - (account payables + accrued expenses)

= 720 - 225

= $495 million.

2) NET WORKING CAPITAL:

= Total current assets - total current liabs.

= 720 - 345

= $375 million.